# Instance-Adaptive Keypoint Learning with Local-to-Global Geometric Aggregation for Category-Level Object Pose Estimation

### v3

@@ Abstract
Category-level object pose estimation aims to predict the 6D pose and size of previously unseen instances from predefined categories, requiring strong generalization across diverse object instances. Although many previous methods attempt to mitigate intra-class variations, they often struggle with instances exhibiting complex geometries or significant deviations from canonical shapes. To address this issue, we propose INKL-Pose, a novel category-level object pose estimation framework that enables INstance-adaptive Keypoint Learning with local-to-global geometric aggregation. Specifically, our method first predicts semantically consistent and geometrically informative keypoints using an Instance-Adaptive Keypoint Detector, then refines them: (1) a Local Keypoint Feature Aggregator capturing fine-grained geometries, and (2) a Global Keypoint Feature Aggregator using bidirectional Mamba for structural consistency. To enable bidirectional modeling in Mamba, we introduce a simple yet effective Feature Sequence Flipping strategy that preserves spatial coherence while constructing backward feature sequence. Additionally, we design a surface loss and a separation loss to encourage uniform coverage and spatial diversity in keypoint distribution. The resulting keypoints are mapped to a canonical space for 6D pose and size regression. Extensive experiments on CAMERA25, REAL275, and HouseCat6D show that INKL-Pose achieves state-of-the-art performance with 16.7M parameters and runs at 36 FPS on an NVIDIA RTX 4090D GPU.
Index Terms:
Object pose estimation, keypoint learning, geometric feature aggregation.
@@ I Introduction
Object pose estimation is a fundamental yet challenging problem in computer vision and robotics, with broad applications in robotic grasping [1], autonomous driving [2], and 3D scene understanding [3]. Existing approaches can be broadly classified into instance-level [4, 5, 6, 7, 8, 9] and category-level [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] methods. Instance-level methods rely on specific CAD models to estimate the poses of known objects. However, their heavy dependence on accurate CAD models makes them difficult to generalize to real-world scenarios, where objects are often novel and lack corresponding models. This severely limits their applicability to unseen instances within the same category or to objects with variations in shape and appearance. As a more flexible alternative, category-level object pose estimation aims to predict the pose and size of previously unseen instances from predefined object categories. Despite its potential, this task remains considerably more challenging due to the significant intra-class variations in object shape, texture, and structure.
@@
Figure 1: Visualization of keypoints in NOCS space predicted by AG-Pose [20] and the proposed INKL-Pose on camera instances with complex geometry. Keypoint colors indicate their NOCS errors, which are linearly mapped within the range $[0,0.2]$ from green (error = 0) to red (error = 0.2). AG-Pose tends to produce clustered and outlier-prone keypoints with higher NOCS errors, whereas INKL-Pose generates spatially diverse and well-distributed keypoints that uniformly cover the object surface with significantly lower errors.
@@
Recent advances in correspondence-based methods [10, 11, 12, 13, 14] have shown promising performance in category-level object pose estimation. However, these approaches typically rely on static priors that restrict their adaptability to novel object instances. To address this limitation, shape-prior-free methods [15, 16, 17, 18, 19, 20] have drawn increasing attention due to their enhanced generalization to novel object instances. Among these methods, AG-Pose [20] proposes a keypoint-based framework that predicts representative keypoints for each instance and establishes keypoint-level correspondences, achieving significant improvements in handling structural discrepancies. Despite its success, AG-Pose derives keypoint features solely from local attention within fixed neighborhoods, combined with a simple global average pooling strategy. This design limits its capacity to capture fine-grained geometric cues and to model global structural consistency. Although AG-Pose incorporates losses to encourage uniformity and separation in keypoint distribution, their effectiveness diminishes when dealing with objects exhibiting complex or irregular geometries. As illustrated in Fig. 1, the resulting keypoints often cluster in limited regions, leading to incomplete surface representation and substantial NOCS errors.
@@
To address the above challenges, we propose INKL-Pose, a novel category-level object pose estimation method that enables instance-adaptive keypoint learning with local-to-global geometric aggregation. Specifically, we first extract and fuse RGB and point cloud features using a Pointwise Feature Encoder, and then generate a sparse set of semantically consistent and geometrically informative keypoints through an Instance-Adaptive Keypoint Detector. These keypoints are subsequently refined by two complementary modules: (1) a Local Keypoint Feature Aggregator that captures fine-grained geometric details from spatial neighborhoods, and (2) a Global Keypoint Feature Aggregator that utilizes a bidirectional Mamba architecture to model long-range dependencies while maintaining structural consistency across the object. To handle the unordered and non-causal nature of point clouds, we introduce a simple yet effective Feature Sequence Flipping strategy that facilitates bidirectional perception in the feature space while preserving the original spatial structure. Additionally, we design a surface loss and a separation loss to explicitly encourage uniform coverage and spatial diversity among keypoints. The refined keypoints are then mapped to a canonical space to regress the object’s 6D pose and size. As illustrated in Fig. 1, even on complex camera instances, the keypoints predicted by our INKL-Pose still achieve accurate and consistent surface coverage, and exhibit significantly lower NOCS errors compared to those predicted by AG-Pose [20].
@@
  • Extensive experiments on CAMERA25, REAL275, and HouseCat6D demonstrate that INKL-Pose achieves state-of-the-art performance on both synthetic and real-world benchmarks, significantly outperforming prior methods.
@@ II-A Instance-Level Object Pose Estimation
Instance-level object pose estimation aims to infer the precise pose of an object based on its corresponding 3D CAD model. Existing approaches in this domain can be broadly divided into three categories: direct regression methods [4, 5], direct voting methods [6, 7] and dense 2D-3D correspondence methods [8, 9]. Direct regression methods utilize end-to-end networks to directly predict the object’s rotation and translation parameters from extracted visual features. Although this approach simplifies the pose estimation pipeline, it suffers from reduced accuracy in cluttered or occluded scenes due to the limitations of direct mapping. Direct voting methods, on the other hand, estimate pose candidates and their corresponding confidence scores at the pixel or point level, ultimately selecting the pose with the highest confidence. While these methods are capable of capturing detailed local information and generally yield strong performance, the dense voting process can be computationally intensive and time-consuming. Alternatively, the third category of instance-level methods follows a correspondence-based paradigm, which first establishes dense 2D–3D correspondences between RGB images and their associated 3D CAD models, and then estimates the object pose using algorithms such as Perspective-n-Point (PnP) or least-squares optimization. Unlike direct regression or voting-based approaches, this method explicitly decouples feature matching from pose estimation, enabling more reliable, geometry-informed inference when accurate correspondences are available. However, their performance heavily relies on the availability of precise CAD models and robust feature descriptors, both of which are often difficult to obtain in real-world scenarios.
@@ II-B Category-Level Object Pose Estimation
Unlike instance-level approaches that rely on known CAD models, category-level 6D object pose estimation aims to estimate the pose and size of previously unseen instances from predefined object categories. This setting is inherently more general and challenging due to substantial intra-class variations among object instances. To address these challenges, NOCS [10] introduces the Normalized Object Coordinate Space, which maps different instances to a unified canonical frame, thereby mitigating variations in scale and geometry. The Umeyama algorithm [22] is then applied to recover the object pose via similarity transformation. To better handle intra-class variations, SPD [11] proposes a deformation network that reconstructs object geometry by learning instance-specific deformations from category-level shape priors, followed by dense correspondence matching between the reconstructed canonical model and the observed point cloud. This paradigm has inspired a series of follow-up works [12, 13, 14], which further explore the extraction of more robust features and the development of more efficient deformation strategies. However, the static nature of shape priors during both training and inference limits adaptability and hinders generalization to objects with substantial structural variations.
To overcome this limitation, Query6DoF [15] introduces class-specific learnable queries that are jointly optimized with the pose estimation network, enabling the shape priors to adapt dynamically during training. In addition, shape-prior-free methods [16, 17, 18, 19, 20] have attracted increasing attention for their improved generalization to diverse and unseen object geometries. IST-Net [17] adopts an implicit spatial transformation framework that directly learns correspondences between camera-space and world-space features without relying on explicit priors. CLIPose [16] leverages a vision-language model to extract rich semantic features from both text and images, aligning them with point cloud representations via contrastive learning. SecondPose [19] integrates geometric features with semantic embeddings from DINOv2 [23], supervised under SE(3)-consistent constraints to improve cross-instance generalization. SpherePose [21] employs a shared spherical proxy representation to address semantic inconsistency arising from shape-dependent canonical spaces. CleanPose [24] integrates causal learning with knowledge distillation to address the spurious correlations introduced by confounding factors in pose estimation. More recently, AG-Pose [20] proposes detecting instance-adaptive keypoints and establishing keypoint-level correspondences to enhance robustness against structural variations across object instances, achieving notable performance improvements.
@@ II-C State Space Models
State Space Models (SSMs) [25], originating from the classical Kalman filter, have attracted increasing interest for their ability to model long-range dependencies with linear computational complexity, while also supporting parallel training. Variants such as S4 [25] introduce diagonal state matrices to reduce memory and computational overhead. Building upon this foundation, Mamba [26] further enhances the SSM framework by incorporating a selective mechanism that compresses context information based on input-dependent parameterization, improving both efficiency and performance. Since the original Mamba operates in a causal manner, where each output token can only attend to preceding tokens, subsequent work such as Vision Mamba [27] introduces bidirectional Mamba to enable information flow in both directions and improve global interaction.
Motivated by these advances, recent studies have adapted the Mamba architecture for 3D point cloud analysis. For instance, PointMamba [28] transforms unordered point cloud data into ordered 1D sequences using Hilbert and Trans-Hilbert curves, preserving spatial continuity and semantic relevance. PCM [29] further utilizes bidirectional Mamba with a point sequence flipping strategy to allow each point to access features from any other point in the cloud. Inspired by the success of Mamba in long-range sequence modeling, we explore its potential for category-level object pose estimation. In particular, we leverage bidirectional Mamba to model globally consistent geometric structures across varied object instances, and propose a novel bidirectional input sequence generation strategy tailored to the unordered and non-causal nature of point cloud data.
@@ III-B Pointwise Feature Encoder
Given an RGB image $\mathcal{I}_{o}$ and its corresponding point cloud $\mathcal{P}_{o}$, the Pointwise Feature Encoder (PFE) extracts informative features from both modalities. Specifically, semantic features $\mathcal{F}_{\mathcal{I}}\in\mathbb{R}^{N\times d_{1}}$ are extracted from the RGB image using DINOv2 [23], while geometric features $\mathcal{F}_{\mathcal{P}}\in\mathbb{R}^{N\times d_{2}}$ are obtained from the point cloud via PointNet++ [30]. To emphasize relative geometric relationships and reduce the influence of absolute spatial positions, the point cloud is preprocessed by subtracting the mean of all coordinates. Although this operation benefits relative geometry modeling, it reduces the network’s sensitivity to absolute positional cues, which are essential for accurate pose estimation. To compensate for this loss, a Multi-Layer Perceptron (MLP) layer is employed to encode the original spatial coordinates into a latent positional embedding $\mathcal{E}_{\mathcal{P}}\in\mathbb{R}^{N\times d_{3}}$. Finally, the semantic features $\mathcal{F}_{\mathcal{I}}$, geometric features $\mathcal{F}_{\mathcal{P}}$, and positional embedding $\mathcal{E}_{\mathcal{P}}$ are concatenated and passed through an additional MLP layer for cross-modal feature fusion, yielding the final pointwise feature representation $\mathcal{F}_{o}\in\mathbb{R}^{N\times d}$, which serves as input for subsequent network modules.
@@ III-C Instance-Adaptive Keypoint Detector
Given the challenges of noise and incomplete observations in cluttered environments, we adopt the Instance-Adaptive Keypoint Detector (IAKD) from AG-Pose [20] as our keypoint generation backbone. This detector formulates keypoint generation as a learnable, query-based process that dynamically generates keypoints from the most informative regions of each object. Specifically, we initialize a learnable keypoint query embedding $\mathcal{Q}_{kpt}$ set to zero, providing a neutral starting point that avoids bias toward specific patterns. Through multi-head cross-attention and self-attention mechanisms, these query embeddings iteratively interact with the input features $\mathcal{F}_{o}$ and themselves, progressively refining to capture instance-specific keypoint features. The keypoint probability distribution matrix $\mathcal{W}$ is computed using cosine similarity followed by a softmax function:
@@
[…]
@@
where $\text{NN}_{j}(\mathcal{P}_{kpt}^{i})$ denotes the $j$-th nearest neighbor of the $i$-th keypoint, and $M$ is the number of neighbors considered. In our experiments, we set $M=2$, which provides a lightweight yet effective constraint to mitigate keypoint clustering while maintaining flexibility in spatial layout. The denominator is lower-bounded by a small constant $\epsilon$ = 0.01 to ensure numerical stability.

### v4

@@ Abstract
Category-level object pose estimation aims to predict the 6D pose and size of previously unseen instances from predefined categories, requiring strong generalization across diverse object instances. Although many previous methods attempt to mitigate intra-class variations, they often struggle with instances (i.e., cameras) exhibiting complex geometries or significant deviations from canonical shapes. To address this issue, we propose INKL-Pose, a novel category-level object pose estimation framework that enables INstance-adaptive Keypoint Learning with local-to-global geometric aggregation. Specifically, our method first predicts semantically consistent and geometrically informative keypoints using an Instance-Adaptive Keypoint Detector, then refines them: (1) a Local Keypoint Feature Aggregator capturing fine-grained geometries, and (2) a Global Keypoint Feature Aggregator using bidirectional Mamba for structural consistency. To enable bidirectional modeling in Mamba, we introduce a simple yet effective Feature Sequence Flipping strategy that preserves spatial coherence while constructing backward feature sequence. Additionally, we design a surface loss and a separation loss to encourage uniform coverage and spatial diversity in keypoint distribution. The resulting keypoints are mapped to a canonical space for 6D pose and size regression. Extensive experiments on CAMERA25, REAL275, and HouseCat6D show that INKL-Pose achieves state-of-the-art performance with 16.7M parameters and runs at 36 FPS on an NVIDIA RTX 4090D GPU.
@@ I Introduction
Object pose estimation is a fundamental yet challenging problem in computer vision and robotics, with broad applications in robotic grasping [wang2024oa, liu2024survey], autonomous driving [qian20223d], and 3D scene understanding [wu2024fairscene]. Existing approaches can be broadly classified into instance-level [di2021so, wang2019densefusion, tian2020robust] and category-level [wang2019normalized, lin2024instance, ren2025learning] methods. Instance-level methods rely on specific CAD models to estimate the poses of known objects. However, their heavy dependence on accurate CAD models makes them difficult to generalize to real-world scenarios, where objects are often novel and lack corresponding models. This severely limits their applicability to unseen instances within the same category or to objects with variations in shape and appearance. As a more flexible alternative, category-level object pose estimation aims to predict the pose and size of previously unseen instances from predefined object categories. Despite its potential, this task remains considerably more challenging due to the significant intra-class variations in object shape, texture, and structure.
@@
Figure 1: Visualization of keypoints in NOCS space predicted by AG-Pose [lin2024instance] and the proposed INKL-Pose on camera instances with complex geometry. Keypoint colors indicate their NOCS errors, which are linearly mapped within the range $[0,0.2]$ from green (error = 0) to red (error = 0.2). AG-Pose tends to produce clustered and outlier-prone keypoints with higher NOCS errors, whereas INKL-Pose generates spatially diverse and well-distributed keypoints that uniformly cover the object surface with significantly lower errors.
@@
Recent advances in correspondence-based methods [tian2020shape, lin2022category, zou2024learning] have shown promising performance in category-level object pose estimation. However, these approaches typically rely on static priors that restrict their adaptability to novel object instances. To address this limitation, shape-prior-free methods [lin2023vi, chen2024secondpose, lin2024instance] have drawn increasing attention due to their enhanced generalization to novel object instances. Among these methods, AG-Pose [lin2024instance] proposes a keypoint-based framework that predicts representative keypoints for each instance and establishes keypoint-level correspondences, achieving significant improvements in handling structural discrepancies. Despite its success, AG-Pose derives keypoint features solely from local attention within fixed neighborhoods, combined with a simple global average pooling strategy. This design limits its capacity to capture fine-grained geometric cues and to model global structural consistency. Although AG-Pose incorporates losses to encourage uniformity and separation in keypoint distribution, their effectiveness diminishes when dealing with objects exhibiting complex or irregular geometries. As illustrated in Fig. 1, the resulting keypoints often cluster in limited regions, leading to incomplete surface representation and substantial NOCS errors.
@@
To address the above challenges, we propose INKL-Pose, a novel category-level object pose estimation method that enables instance-adaptive keypoint learning with local-to-global geometric aggregation. Specifically, we first extract and fuse RGB and point cloud features using a Pointwise Feature Encoder, and then generate a sparse set of semantically consistent and geometrically informative keypoints through an Instance-Adaptive Keypoint Detector. These keypoints are subsequently refined by two complementary modules: (1) a Local Keypoint Feature Aggregator that captures fine-grained geometric details from spatial neighborhoods, and (2) a Global Keypoint Feature Aggregator that utilizes a bidirectional Mamba architecture to model long-range dependencies while maintaining structural consistency across the object. To handle the unordered and non-causal nature of point clouds, we introduce a simple yet effective Feature Sequence Flipping strategy that facilitates bidirectional perception in the feature space while preserving the original spatial structure. Additionally, we design a surface loss and a separation loss to explicitly encourage uniform coverage and spatial diversity among keypoints. The refined keypoints are then mapped to a canonical space to regress the object’s 6D pose and size. As illustrated in Fig. 1, even on complex camera instances, the keypoints predicted by our INKL-Pose still achieve accurate and consistent surface coverage, and exhibit significantly lower NOCS errors compared to those predicted by AG-Pose [lin2024instance].
@@ II-A Instance-Level Object Pose Estimation
Instance-level object pose estimation aims to infer the precise pose of an object based on its corresponding 3D CAD model. Existing approaches in this domain can be broadly divided into three categories: direct regression methods [di2021so, wang2019densefusion], direct voting methods [tian2020robust, zhou2021pr] and dense 2D-3D correspondence methods [park2019pix2pose, xu2024rnnpose]. Direct regression methods utilize end-to-end networks to directly predict the object’s rotation and translation parameters from extracted visual features. Although this approach simplifies the pose estimation pipeline, it suffers from reduced accuracy in cluttered or occluded scenes due to the limitations of direct mapping. Direct voting methods, on the other hand, estimate pose candidates and their corresponding confidence scores at the pixel or point level, ultimately selecting the pose with the highest confidence. While these methods are capable of capturing detailed local information and generally yield strong performance, the dense voting process can be computationally intensive and time-consuming. Alternatively, the third category of instance-level methods adopts a correspondence-based paradigm, which first establishes dense 2D–3D correspondences between RGB images and the complete 3D CAD model, and subsequently estimates the object pose using algorithms such as Perspective-n-Point or least-squares optimization. Unlike direct regression or voting-based approaches, this method explicitly decouples feature matching from pose estimation, enabling more reliable, geometry-informed inference when accurate correspondences are available. However, their performance heavily relies on the availability of precise CAD models and robust feature descriptors, both of which are often difficult to obtain in real-world scenarios.
@@ II-B Category-Level Object Pose Estimation
Unlike instance-level approaches that rely on known CAD models, category-level 6D object pose estimation aims to estimate the pose and size of previously unseen instances from predefined object categories. This setting is inherently more general and challenging due to substantial intra-class variations among object instances. To address these challenges, NOCS [wang2019normalized] introduces the Normalized Object Coordinate Space, which maps different instances to a unified canonical frame, thereby mitigating variations in scale and geometry. The Umeyama algorithm [umeyama1991least] is then applied to recover the object pose via similarity transformation. To better handle intra-class variations, SPD [tian2020shape] proposes a deformation network that reconstructs object geometry by learning instance-specific deformations from category-level shape priors, followed by dense correspondence matching between the reconstructed canonical model and the observed point cloud. This paradigm has inspired a series of follow-up works [tian2020shape, lin2022category, zou2024learning], which further explore the extraction of more robust features and the development of more efficient deformation strategies. However, the static nature of shape priors during both training and inference limits adaptability and hinders generalization to objects with substantial structural variations.
To overcome this limitation, Query6DoF [wang2023query6dof] introduces class-specific learnable queries that are jointly optimized with the pose estimation network, enabling the shape priors to adapt dynamically during training. In addition, shape-prior-free methods [liu2023net, lin2024clipose, chen2024secondpose, lin2024instance, lin2025cleanpose] have attracted increasing attention for their improved generalization to diverse and unseen object geometries within the same category. IST-Net [liu2023net] adopts an implicit spatial transformation framework that directly learns correspondences between camera-space and world-space features without relying on explicit priors. CLIPose [lin2024clipose] leverages a vision-language model to extract rich semantic features from both text and images, aligning them with point cloud representations via contrastive learning. SecondPose [chen2024secondpose] integrates geometric features with semantic embeddings from DINOv2 [oquab2023dinov2], supervised under SE(3)-consistent constraints to improve cross-instance generalization. SpherePose [ren2025learning] employs a shared spherical proxy representation to address semantic inconsistency arising from shape-dependent canonical spaces. CleanPose [lin2025cleanpose] integrates causal learning with knowledge distillation to address the spurious correlations introduced by confounding factors in pose estimation. More recently, AG-Pose [lin2024instance] proposes detecting instance-adaptive keypoints and establishing keypoint-level correspondences to enhance robustness against structural variations across object instances, achieving notable performance improvements.
@@ II-C State Space Models
State Space Models (SSMs) [gu2022efficiently], originating from the classical Kalman filter, have attracted increasing interest for their ability to model long-range dependencies with linear computational complexity, while also supporting parallel training. Variants such as S4 [gu2022efficiently] introduce diagonal state matrices to reduce memory and computational overhead. Building upon this foundation, Mamba [gu2024mamba] further enhances the SSM framework by incorporating a selective mechanism that compresses context information based on input-dependent parameterization, improving both efficiency and performance. Since the original Mamba operates in a causal manner, where each output token can only attend to preceding tokens, subsequent work such as Vision Mamba [vim] introduces bidirectional Mamba to enable information flow in both directions and improve global interaction.
Motivated by these advances, recent studies [li2025enhancing, liang2024pointmamba, zhang2025point] have adapted the Mamba architecture for 3D point cloud analysis. For instance, PointMamba [liang2024pointmamba] transforms unordered point cloud data into ordered 1D sequences using Hilbert and Trans-Hilbert curves, preserving spatial continuity and semantic relevance. PCM [zhang2025point] further utilizes bidirectional Mamba with a point sequence flipping strategy to allow each point to access features from any other point in the cloud. Inspired by the success of Mamba in long-range sequence modeling, we explore its potential for category-level object pose estimation. In particular, we leverage bidirectional Mamba to model globally consistent geometric structures across varied object instances, and propose a novel bidirectional input sequence generation strategy tailored to the unordered and non-causal nature of point cloud data.
@@ III-B Pointwise Feature Encoder
Given an RGB image $\mathcal{I}_{o}$ and its corresponding point cloud $\mathcal{P}_{o}$, the Pointwise Feature Encoder (PFE) extracts informative features from both modalities. Specifically, semantic features $\mathcal{F}_{\mathcal{I}}\in\mathbb{R}^{N\times d_{1}}$ are extracted from the RGB image using DINOv2 [oquab2023dinov2], while geometric features $\mathcal{F}_{\mathcal{P}}\in\mathbb{R}^{N\times d_{2}}$ are obtained from the point cloud via PointNet++ [qi2017pointnet++]. To emphasize relative geometric relationships and reduce the influence of absolute spatial positions, the point cloud is preprocessed by subtracting the mean of all coordinates. Although this operation benefits relative geometry modeling, it reduces the network’s sensitivity to absolute positional cues, which are essential for accurate pose estimation. To compensate for this loss, a Multi-Layer Perceptron (MLP) layer is employed to encode the original spatial coordinates into a latent positional embedding $\mathcal{E}_{\mathcal{P}}\in\mathbb{R}^{N\times d_{3}}$. Finally, the semantic features $\mathcal{F}_{\mathcal{I}}$, geometric features $\mathcal{F}_{\mathcal{P}}$, and positional embedding $\mathcal{E}_{\mathcal{P}}$ are concatenated and passed through an additional MLP layer for cross-modal feature fusion, yielding the final pointwise feature representation $\mathcal{F}_{o}\in\mathbb{R}^{N\times d}$, which serves as input for subsequent network modules.
@@ III-C Instance-Adaptive Keypoint Detector
Given the challenges of noise and incomplete observations in cluttered environments, we adopt the Instance-Adaptive Keypoint Detector (IAKD) from AG-Pose [lin2024instance] as our keypoint generation backbone. This detector formulates keypoint generation as a learnable, query-based process that dynamically generates keypoints from the most informative regions of each object. Rather than relying on fixed strategies like Farthest Point Sampling (FPS), which select keypoints solely based on spatial distances and fail to adapt to the geometric diversity across instances, the IAKD learns to predict instance-specific keypoints in a data-driven and context-aware manner. Specifically, we initialize a learnable keypoint query embedding $\mathcal{Q}_{kpt}$ set to zero, providing a neutral starting point that avoids bias toward specific patterns. Through multi-head cross-attention and self-attention mechanisms, these query embeddings iteratively interact with the input features $\mathcal{F}_{o}$ and themselves, progressively refining to capture instance-specific keypoint features. The keypoint probability distribution matrix $\mathcal{W}$ is computed using cosine similarity followed by a softmax function:
@@
$\begin{aligned} \mathcal{L}_{surf}(\mathcal{P}_{fps},\mathcal{P}_{kpt})=&\,\frac{1}{{N_{fps}}}\sum_{x_{i}\in\mathcal{P}_{fps}}\min_{y_{i}\in\mathcal{P}_{kpt}}\|x_{i}-y_{i}\|_{2}^{2}+\\ &\,\frac{1}{{N_{kpt}}}\sum_{y_{i}\in\mathcal{P}_{kpt}}\min_{x_{i}\in\mathcal{P}_{fps}}\|y_{i}-x_{i}\|_{2}^{2}\end{aligned}.$ (3)
@@
where $\text{NN}_{j}(\mathcal{P}_{kpt}^{i})$ denotes the $j$th nearest neighbor of the $i$th keypoint, and $M$ is the number of neighbors considered. In our experiments, we set $M=2$, which provides a lightweight yet effective constraint to mitigate keypoint clustering while maintaining flexibility in spatial layout. The denominator is lower-bounded by a constant $\epsilon$ to ensure numerical stability.